OpenStudy (anonymous):

State the vertical asymptote of the rational function. f(x)=(x-3)(x+4)/x^2-1

OpenStudy (johnweldon1993):

how do we find a vertical asymptote? when the denominator = 0....that makes it so when does x² - 1 = 0?

OpenStudy (anonymous):

when x = 0?

OpenStudy (johnweldon1993):

that's....a start....remember x is squared....so what else can it be?

OpenStudy (johnweldon1993):

for this particular problem....there will be 2 vertical asymptotes...because x² - 1 = 0 has 2 answers

OpenStudy (johnweldon1993):

and you deleted your answer of x = 1....that was a correct answer

OpenStudy (anonymous):

oh, sorry, only because I put w=1, I was going to retype it, but i had to take my dogs out

OpenStudy (anonymous):

Are you there?

OpenStudy (johnweldon1993):

sorry about that...yes.....and yes okay so x = 1 is an answer how else can x² - 1 = 0? we know that 1² - 1 = 0 right? as you stated what else can be ² to make 1?

OpenStudy (anonymous):

-1?

OpenStudy (johnweldon1993):

correct! :) so your 2 answers would be a vertical asymptote at 1, and -1

OpenStudy (anonymous):

Oh, And that's it?

OpenStudy (anonymous):

Thanks:)

OpenStudy (johnweldon1993):

that's all it's asking for is to find the vertical asymptotes....so that would be it :)